leila1234567:

u=12ba+9a solve for a

umm:

Given \Let's flip the equation giving us\[12ab+9a=u\]Now we get to factor our variable a\[a(12b+9)=u\]Divide both sides\[\frac{ a(12b+9) }{ 12b+9 }=\frac{ u }{ 12b+9 }\]Our final answer is \[a= \frac{ u }{ 12b+9 }\]

leila1234567:

ok thank you

umm:

No problem

llamaleah:

h